OpenStudy (anonymous):

Can someone separate this differential equation for me? (dy/dx)=2xy+4x

OpenStudy (freckles):

2xy+4x=2x(y+2) dy/dx=2x(y+2) divide both sides by (y+2) and multiply both sides by dx then integrate both sides

OpenStudy (anonymous):

dy/dx = 2x(y+2) separate variables(crossmultiply) dy/(y+2) = 2x dx Integrate : ln(y+2) = x^2 + C

OpenStudy (anonymous):